OpenStudy (anonymous):

If f(x) = x2 + 1, find f(3).

OpenStudy (anonymous):

plug value of 3 in x;

OpenStudy (anonymous):

\[f(x) = x^2 + 1\] \[f(3) = 3^2 + 1\] \[f(3) = 9 + 1\] \[f(3) = 10\]